# Multi-Domain Graph Foundation Models: Robust Knowledge Transfer via Topology Alignment

### v1

@@ Abstract
Recent advances in CV and NLP have inspired researchers to develop general-purpose graph foundation models through pre-training across diverse domains. However, a fundamental challenge arises from the substantial differences in graph topologies across domains. Additionally, real-world graphs are often sparse and prone to noisy connections and adversarial attacks. To address these issues, we propose the Multi-Domain Graph Foundation Model (MDGFM), a unified framework that aligns and leverages cross-domain topological information to facilitate robust knowledge transfer. MDGFM bridges different domains by adaptively balancing features and topology while refining original graphs to eliminate noise and align topological structures. To further enhance knowledge transfer, we introduce an efficient prompt-tuning approach. By aligning topologies, MDGFM not only improves multi-domain pre-training but also enables robust knowledge transfer to unseen domains. Theoretical analyses provide guarantees of MDGFM’s effectiveness and domain generalization capabilities. Extensive experiments on both homophilic and heterophilic graph datasets validate the robustness and efficacy of our method.
@@ 1 Introduction
Despite significant progress, a substantial gap remains in fully understanding the richness and diversity of graph structural knowledge. Existing methods primarily rely on fixed graph topologies and apply uniform encoding mechanisms across all domains (Zhao et al., 2024; Yu et al., 2025), which severely limits their generalizability across diverse domains (Zhang et al., 2024a). In this paper, we revisit the multi-domain graph foundation model from a structural perspective and address a critical challenge: topology alignment across different domains. Achieving effective topology alignment is a non-trivial task, presenting two primary challenges that must be overcome to advance toward a truly general-purpose graph model.
First, structural knowledge across domains often exhibits significant semantic differences. For instance, citation networks predominantly display highly homophilic patterns, whereas social networks and webpage link graphs frequently contain numerous heterophilic edges (Zheng et al., 2022). These differences necessitate domain-specific adaptations of encoding mechanisms. As a result, there is an urgent need for a unified framework that can adaptively capture critical information from both features and topology while effectively learning domain-invariant knowledge. Such a framework is essential to ensure robust generalization to downstream domains (Li et al., 2022).
@@ 2.2 Multi-domain Generalization
Recently, domain generalization on graphs has gained attention. These methods integrate and extract knowledge from multiple source domains during the upstream pre-training phase, enabling the transfer of this knowledge to tackle various graph-related tasks in previously unseen downstream domains. For example, GCOPE (Zhao et al., 2024) integrates multi-source graph topologies during the pre-training phase by introducing interconnected virtual nodes. Additionally, MDGPT (Yu et al., 2025) incorporates domain-specific tokens in the pre-training phase to align node features from different domains, and employs prompt-tuning during downstream tasks for efficient knowledge transfer. Despite these advancements, significant gaps remain in understanding the semantic differences and reliability of multi-domain topologies. Consequently, there is an urgent need to design a universal foundation model capable of achieving robust and generalized knowledge transfer across domains.
@@ 4.2 Graph Topology-aware Alignment
Finally, consistent with existing GSL methods (Li et al., 2024), we apply post-processing techniques to reconstruct the refined adjacency matrix. Specifically, a similarity matrix is first computed using $H_{i}$ and then sparsified via $k$-nearest neighbors ($k$NN). In practice, we employ $k$NN sparsification with its locality-sensitive approximation to enhance efficiency (Fatemi et al., 2021). Subsequently, operations such as Symmetrization, Activation, and Normalization are performed sequentially to produce the final $A^{\prime}_{i}$. Detailed implementation steps are provided in Appendix E. Building on this topology-aware refinement process, we propose a general pretraining framework capable of generating high-quality embeddings across multiple domains using contrastive learning. By optimizing a contrastive-based objective, the model learns generalized representations by leveraging rich sample-to-sample relationships from diverse perspectives (Yao et al., 2022). Mathematically, we maximize the mutual information between $G_{i1}=(A_{i},X_{i}^{\prime})$ and $G_{i2}=(A_{i}^{\prime},X_{i}^{\prime})$ as follows:
@@
where we use $\dagger$ to indicate positive samples when computing similarities, often utilizing the identity matrix $I_{e}$ for this purpose. The second term of $\mathcal{L}$ incorporates the refined graph structure $A_{i}^{\prime}$ to increase the number of positive samples. Detailed computations of the loss function are provided in Appendix F. This loss formulation reduces the information gap between the original and refined graphs, ensuring that global structural properties are preserved during optimization. By maintaining semantic consistency in the graph structure, it enhances alignment and adaptability for downstream tasks.
@@ 4.3 Knowledge Transfer to Downstream Domain
Similar to in-context learning, downstream adaptation aims to enhance a model’s ability to learn tasks using only a few examples provided as demonstrations (Dong et al., 2022), highlighting the need for upstream knowledge transfer. Inspired by the “pre-training &\&& prompting” paradigm, our framework employs a dual-prompt strategy. Specifically, the meta prompt $p_{m}$ focuses on adjusting the distribution of learned knowledge by modeling the relationships between the target and source domains. This can be formalized as a function defined for feature $X$:
@@ 5 Theoretical Analysis
To minimize the upper error bound, we introduce the concept of invariant graph learning:
@@ 6.1 Experimental Setups
Datasets. To ensure a comprehensive comparison, we conduct experiments on six primary datasets, including three homophilic graphs—Cora (Sen et al., 2008), Citeseer (Sen et al., 2008), and Pubmed (Namata et al., 2012)—and three heterophilic graphs—Cornell, Chameleon, and Squirrel (Pei et al., 2020). Additionally, we include Penn94 (Traud et al., 2012), a large-scale graph dataset, as a downstream target domain. Detailed statistical information on these datasets is provided in Appendix B.
@@
For a fair comparison, we use GCN as the backbone for all methods. To evaluate performance on unseen target domains, we focus on scenarios where the model generalizes to domains not encountered during pre-training. For the Cora, Citeseer, Pubmed, Chameleon, Squirrel, and Cornell datasets, we designate one dataset as the downstream target domain while using the remaining five as source domains during pre-training. Additionally, we extend this setup by using all six datasets as source domains and applying Penn94 as the target domain for few-shot learning. More details are provided in Appendix D.
@@ 6.4 Domain Sensitivity (RQ3)
Removed Data GCOPE MDGPT MDGFM
@@ 6.5 Robustness Analysis (RQ4)
Furthermore, we employ Metattack (Zügner & Günnemann, 2019), an advanced adversarial attack that perturbs the training data to degrade the model performance after training. Under identical attack conditions, we compare MDGFM against multi-domain methods. As shown in Figure 3, the performance of GCOPE and MDGPT deteriorates as attack intensity increases. Notably, our model is quite stable and consistently outperforms all baselines across all attack scenarios, which can be attributed to our proposed topology alignment. Further analysis, including a sensitivity study of hyperparameters, is provided in Appendix C.
@@ Appendix A Notations
Appendix B Datasets
@@
Appendix C Additional Experiments
C.1 Expansion of Robustness Experiments
@@ C.2 Sensitivity Analysis
Appendix D Experimental Details
@@
Appendix E Methodology Details
@@
Appendix F Details of Loss Functions

### v2

@@ Abstract
Recent advances in CV and NLP have inspired researchers to develop general-purpose graph foundation models through pre-training across diverse domains. However, a fundamental challenge arises from the substantial differences in graph topologies across domains. Additionally, real-world graphs are often sparse and prone to noisy connections and adversarial attacks. To address these issues, we propose the Multi-Domain Graph Foundation Model (MDGFM), a unified framework that aligns and leverages cross-domain topological information to facilitate effective and robust knowledge transfer. MDGFM bridges different domains by adaptively balancing features and topology while refining original graphs to eliminate noise and align topological structures. To further enhance knowledge transfer, we introduce an efficient prompt-tuning approach. By aligning topologies, MDGFM not only improves multi-domain pre-training but also enables robust knowledge transfer to unseen domains. Theoretical analyses provide guarantees of MDGFM’s effectiveness and domain generalization capabilities. Extensive experiments on both homophilic and heterophilic graph datasets validate the robustness and efficacy of our method. Our code is available at https://github.com/wbkzwqtzw/MDGFM.
@@ 1 Introduction
Despite significant progress, a substantial gap remains in fully understanding the richness and diversity of graph structural knowledge. Existing methods primarily rely on fixed graph topologies and apply uniform encoding mechanisms across all domains (Zhao et al., 2024; Yu et al., 2025), which severely limits their generalizability across diverse domains (Zhang et al., 2024). In this paper, we revisit the multi-domain graph foundation model from a structural perspective and address a critical challenge: topology alignment across different domains. Achieving effective topology alignment is a non-trivial task, presenting two primary challenges that must be overcome to advance toward a truly general-purpose graph model.
First, structural knowledge across domains often exhibits significant semantic differences. For instance, citation networks predominantly display highly homophilic patterns, whereas social networks and webpage link graphs frequently contain numerous heterophilic edges (Zheng et al., 2022; Shen and Kang, 2025; Xie et al., 2025). These differences necessitate domain-specific adaptations of encoding mechanisms (Li et al., 2024; Pan and Kang, 2023). As a result, there is an urgent need for a unified framework that can adaptively capture critical information from both features and topology while effectively learning domain-invariant knowledge. Such a framework is essential to ensure robust generalization to downstream domains (Li et al., 2022; Xie et al., 2025).
@@ 2.2 Multi-domain Generalization
Recently, domain generalization on graphs has gained attention (Wang et al., 2024). These methods integrate and extract knowledge from multiple source domains during the upstream pre-training phase, enabling the transfer of this knowledge to tackle various graph-related tasks in previously unseen downstream domains (Fang et al., 2025). For example, GCOPE (Zhao et al., 2024) integrates multi-source graph topologies during the pre-training phase by introducing interconnected virtual nodes. Additionally, MDGPT (Yu et al., 2025) incorporates domain-specific tokens in the pre-training phase to align node features from different domains, and employs prompt-tuning during downstream tasks for efficient knowledge transfer. Despite these advancements, significant gaps remain in understanding the semantic differences and reliability of multi-domain topologies. Consequently, there is an urgent need to design a universal foundation model capable of achieving robust and generalized knowledge transfer across domains.
@@ 4.2 Graph Topology-aware Alignment
Finally, consistent with existing GSL methods (Li et al., 2024), we apply post-processing techniques to reconstruct the refined adjacency matrix. Specifically, a similarity matrix is first computed using $H_{i}$ and then sparsified via $k$-nearest neighbors ($k$NN). In practice, we employ $k$NN sparsification with its locality-sensitive approximation to enhance efficiency (Fatemi et al., 2021). Subsequently, operations such as Symmetrization, Activation, and Normalization are performed sequentially to produce the final $A^{\prime}_{i}$. Detailed implementation steps are provided in Appendix F. Building on this topology-aware refinement process, we propose a general pretraining framework capable of generating high-quality embeddings across multiple domains using contrastive learning. By optimizing a contrastive-based objective, the model learns generalized representations by leveraging rich sample-to-sample relationships from diverse perspectives (Yao et al., 2022). Mathematically, we maximize the mutual information between $G_{i1}=(A_{i},X_{i}^{\prime})$ and $G_{i2}=(A_{i}^{\prime},X_{i}^{\prime})$ as follows:
@@
where we use $\dagger$ to indicate positive samples when computing similarities, often utilizing the identity matrix $I_{e}$ for this purpose. The second term of $\mathcal{L}$ incorporates the refined graph structure $A_{i}^{\prime}$ to increase the number of positive samples. Detailed computations of the loss function are provided in Appendix G. This loss formulation reduces the information gap between the original and refined graphs, ensuring that global structural properties are preserved during optimization. By maintaining semantic consistency in the graph structure, it enhances alignment and adaptability for downstream tasks.
@@ 4.3 Knowledge Transfer to Downstream Domain
Similar to in-context learning, downstream adaptation aims to enhance a model’s ability to learn tasks using only a few examples provided as demonstrations (Dong et al., 2024), highlighting the need for upstream knowledge transfer. Inspired by the “pre-training &\&& prompting” paradigm, our framework employs a dual-prompt strategy. Specifically, the meta prompt $p_{m}$ focuses on adjusting the distribution of learned knowledge by modeling the relationships between the target and source domains. This can be formalized as a function defined for feature $X$:
@@ 5 Theoretical Analysis
To minimize the upper error bound, we introduce the concept of invariant graph learning, which is highlighted in our work:
@@ 6.1 Experimental Setups
Datasets. To ensure a comprehensive comparison, we conduct experiments on six primary datasets, including three homophilic graphs—Cora (Sen et al., 2008), Citeseer (Sen et al., 2008), and Pubmed (Namata et al., 2012)—and three heterophilic graphs—Cornell, Chameleon, and Squirrel (Pei et al., 2020). Additionally, we include Penn94 (Traud et al., 2012), a large-scale graph dataset, as a downstream target domain. Detailed statistical information on these datasets is provided in Appendix C.
@@
For a fair comparison, we use GCN as the backbone for all methods. To evaluate performance on unseen target domains, we focus on scenarios where the model generalizes to domains not encountered during pre-training. For the Cora, Citeseer, Pubmed, Chameleon, Squirrel, and Cornell datasets, we designate one dataset as the downstream target domain while using the remaining five as source domains during pre-training. Additionally, we extend this setup by using all six datasets as source domains and applying Penn94 as the target domain for few-shot learning. More details are provided in Appendix E.
@@ 6.4 Domain Sensitivity (RQ3)
Removed Data GCOPE SAMGPT MDGFM
@@ 6.5 Robustness Analysis (RQ4)
Furthermore, we employ Metattack (Zügner & Günnemann, 2019), an advanced adversarial attack that perturbs the training data to degrade the model performance after training. Under identical attack conditions, we compare MDGFM against multi-domain methods. As shown in Figure 3, the performance of GCOPE and MDGPT deteriorates as attack intensity increases. Notably, our model is quite stable and consistently outperforms all baselines across all attack scenarios, which can be attributed to our proposed topology alignment. Further analysis, including a sensitivity study of hyperparameters, is provided in Appendix D. Time complexity analysis is also summarized in Appendix B to show the efficiency of MDGFM.
@@ 7 Conclusion
Impact Statement
This paper aims to advance the field of Machine Learning. While our work has potential societal implications, we do not identify any specific concerns that require particular emphasis at this stage.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62276053).
@@ Appendix A Notations
Appendix B Time complexity analysis
In the pre-training phase, each of the $N$ source graphs is processed independently. For a graph with $|V|$ nodes and $|E|$ edges (here we select $|V|=max\{V_{n}\},|E|=max\{E_{m}\},n,m\in[1,N]$), the model first aligns node features via truncated PCA, which reduces the input dimension from $d^{\prime}$ to $d$ at a cost of $\mathcal{O}(|V|\cdot d\cdot\log d^{\prime})$. As for token lightweight element-wise multiplication, the time complexity is $|V|d$. It then applies locality-sensitive hashing $k$NN for graph structure learning. Denote the batch size of sparse $k$NN as $B$, each requiring $\mathcal{O}(d)$ operations, resulting in $\mathcal{O}(|V|\cdot B\cdot d)$ time for reconstructing the graph. Next, an $L$-layer GCN operates on the refined structure , contributing an additional $\mathcal{O}(L\cdot|V|\cdot d^{2}+L\cdot|E|\cdot d+|V|d)$. Therefore, the total pre-training complexity across all source graphs is $\mathcal{O}\left(N\cdot\left[|V|\cdot d\cdot\log d^{\prime}+|V|\cdot B\cdot d+L\cdot|V|\cdot d^{2}+L\cdot|E|\cdot d\right]\right)$.
Similarly, in the downstream phase, PCA procedure takes $\mathcal{O}(|V_{T}|\cdot d\cdot logd^{\prime})$ time. Prompt fusion and token modulation cost $\mathcal{O}(|V_{T}|\cdot d)$. GSL again uses local sensitive $k$NN, which adds $\mathcal{O}(|V_{T}|\cdot B\cdot d)$. The GCN encoder then performs $L$-layer message passing with cost $\mathcal{O}(L\cdot|V_{T}|\cdot d^{2}+L\cdot|E_{T}|\cdot d+|V_{T}|d)$. Finally, classification is done via prototype matching, where each node compares with $C$ class centroids, yielding $\mathcal{O}(|V_{T}|\cdot C\cdot d)$. Summing these terms, the overall downstream complexity is $\mathcal{O}(|V_{T}|\cdot d\cdot\log d^{\prime}+|V_{T}|\cdot B\cdot d+L\cdot|E_{T}|\cdot d+L\cdot|V_{T}|\cdot d^{2}+|V_{T}|\cdot C\cdot d)$.
Overall, the model scales linearly with the number of nodes and edges, and benefits from efficient structure refinement and modular design like local sensitive $k$NN. We will include this complexity analysis in the final version for completeness.
Appendix C Datasets
@@
Appendix D Additional Experiments
D.1 Expansion of Robustness Experiments
@@ D.2 Sensitivity Analysis
Appendix E Experimental Details
@@
Appendix F Methodology Details
@@
Appendix G Details of Loss Functions